When planning for the big Statistics Department Super Bowl party, it is sometimes not clear what...

  1. When planning for the big Statistics Department Super Bowl party, it is sometimes not clear what kinds of chips that people will want to have available, nor is it clear how many people will actually show up at the party. Because of this Professor Mean takes a 4-sided die to the grocery store, starts at one end of the chips aisle, and travels to the other end. At each different kind of chips, Dr. Mean rolls the die. If it comes up a 4, she purchases the chips for the party. There are 40 different kinds of chips in the aisle. Define random variable x = number of types of chips purchased out of 40.

    1. a) What is the mean of the random variable x?

    2. b) What is the standard deviation of the random variable x?
Answer #1

here as probability of 4 on a dice p=1/6

X is binomial distribution with parameter n=40 and p=1/6

a) mean of the random variable x =np=40*1/6=6.6667

b) standard deviation of the random variable x =sqrt(np(1-p))=sqrt(40*1/6*5/6)=2.357